# Multimodal ”Puppeteer”: An Exploration of Robot Teleoperation Via Virtual Counterpart with LLM-Driven Voice and Gesture Interaction in Augmented Reality

### v1

@@ IV-D Measurements
To evaluate the impact of the multimodal interaction, we employed a combination of performance metrics and UX and usability metrics. Performance was measured using four key indicators: number of cubes, unsuccessful attempts, task completion time (TCT, the maximum is 300s), and number of interruptions, providing insights into task efficiency, accuracy, and user control. For UX and usability assessment, we selected UEQ-S (pragmatic quality and hedonic quality, in [-3, +3 Likert scale], with each consisting of four metrics) [50, 51, 52], a shortened version of UEQ, to balance comprehensive evaluation with experiment duration constraints. Additionally, NASA TLX was utilized to measure the perceived workload in 7-point Likert scale [53, 54]. To further explore usability aspects, we included six customized questions inspired by previous works [55, 56, 52], focusing on system usability and user perceptions, which were also measured using 7-point Likert scale for unifying the process as well as better quantifying the evaluation. The complete hierarchy and definitions of self-developed metrics are detailed in Tab. I. We implemented normality test of each metric, then applied the corresponding statistical analysis.
V Results
V-A Quantitative Analysis
V-A1 Performance Metrics
@@
Number of Cubes. This metric was significantly higher in the GO condition ($M=4.6,SD=2.19$) compared to VG ($M=3.6,SD=1.79,p<0.05$). The result indicates that incorporating voice commands may have introduced additional cognitive load or inefficiencies, leading to a lower number of successfully moved cubes. The observed variability in performance as shown in Fig. 6 suggests differential adaptation to multimodal interaction among participants.
@@
Unsuccessful Attempts. As shown in Fig. 6, no significant difference was found between the two conditions. The mean number of failed attempts was slightly higher in VG ($M=7.57,SD=2.41$) than in GO ($M=7.13,SD=3.06$), but this difference was not statistically significant. This implies that error rates were likely influenced by factors such as individual skill levels or task complexity, rather than the interaction modality.
@@
TCT. Given the 300s’ time limit, most participants in VG reached the maximal time, whereas the GO group had a lower TCT ($M=261s$), achieving successful completion more frequently ($p<0.01$) (Fig. 6). This finding indicates that gesture-only interaction facilitated more efficient task execution, likely due to fewer errors and quicker responses compared to the multimodal VG condition.
@@
Interruptions. This metric slightly more frequent in the GO condition ($M=5.95,SD=4.30$) than in VG (Fig. 6). However, this difference was not statistically significant. The presence of outliers in GO suggests that some users encountered more issues without voice commands, hinting at a possible but inconsistent advantage of multimodal input in reducing interaction errors.
@@
V-A2 UX and Usability Metrics
UEQ-S. An item-wise analysis of UEQ-S revealed that GO tended to exceed the values across most variables compared to VG (Fig. 7). Despite considerable overlap in distributions, certain items – such as Boring-Exciting – showed clearer separation. Notably, the Inefficient–Efficient scale favored GO significantly (p ¡ 0.01).
@@ V-A2 UX and Usability Metrics
For conventional analysis, pragmatic quality revealed a highly significant difference (p ¡ 0.001), with GO ($M=5.19,SD=1.52$) rated higher than VG ($M=4.76,SD=1.64$) (Fig. 7). This suggests that gesture-only interaction was perceived as more efficient, predictable, and user-friendly. Gestures, being direct and immediate, likely contributed to this perception, whereas voice commands may have introduced recognition delays or increased cognitive load. However, hedonic quality did not differ significantly between conditions ($p>0.05$), with GO ($M=5.56,SD=1.10$) and VG ($M=5.47,SD=1.14$), which were rated similarly (Fig. 7). This suggests both modalities were comparably engaging and enjoyable. The multimodal novelty of VG did not significantly enhance user enjoyment, implying that gesture-based input alone is sufficiently intuitive and immersive for such experiences. The overall quality significantly favored GO ($M=5.38,SD=1.34$) over VG ($M=5.12,SD=1.46$) ($p<0.001$) (Fig. 7). This reflects the strong impact of GO’s pragmatic advantages, as hedonic differences were negligible. The results suggest that while voice input may offer contextual benefits, it does not inherently improve UX when combined with gestures. On the contrary, it may introduce complexity and inefficiencies that detract from perceived usability.
@@
NASA TLX. Interestingly as Fig. 7 shows, the VG condition recorded higher scores across most workload categories, contrary to expectations that multimodality would trigger lower scores indicating reduced workload. Notably, VG exhibited more compressed distributions in certain dimensions, pointing to more consistent user perceptions. Significant differences were found in physical demand ($p<0.005$), with GO perceived as more physically demanding. This suggests that integrating voice commands alleviates some physical effort. For Perceived Performance, VG was rated significantly higher than GO ($p<0.005$), implying that participants felt more successful in task execution when only voice interaction was available. Although mental demand showed no significant difference ($p>0.05$), the small effect size ($Cohen’sd=0.26$) indicates a slight increase in mental demand for GO. No significances were observed for Temporal Demand, Effort, and Frustration, suggesting similar perceptions of time pressure, exertion, and frustration across both conditions.
Customized Metrics. Compared to standard instruments, customized metrics displayed greater variability (Fig. 7). Some GO metrics had higher medians but wider spread, while VG values tended to be lower and more consistent, indicating possible usability concerns. Significant differences were observed in perceived effectiveness ($p<0.005$), perceived intent to use ($p<0.005$), and satisfaction ($p<0.005$), with GO outperforming VG in all three. This suggests that participants found gesture-only interaction more effective, easier to use, and more satisfying overall. However, the broad variance indicates individual user preferences played a substantial role, especially in other customized metrics not reaching significant differences.
V-B Qualitative Analysis
We now present the qualitative findings derived from the post-study interviews to gain deeper insight into participants’ experiences and perceptions, obtained from a thematic analysis. One author led the coding process and theme development using an affinity diagram approach, facilitated through a shared Miro board for collaborative access. Two additional authors refined and revised the emerging themes with the assistance of ChatGPT. Finally, the themes were finalized through joint discussion among all authors. The main themes with hierarchical details are distilled in Fig. 8.
@@ V-B Qualitative Analysis
V-B1 Overall Experience
Positive feedback on intuitiveness & novelty. Participants found the system novel and engaging, appreciating its innovative interaction methods. One participant stated, “I feel like it’s a very novel experience. You can interact with virtual stuff, like the robot model, as well as controlling the real robot. It’s very interesting.” Another noted, “the whole experience is quite nice, everything is very smooth,” underlining the potential for intuitive interactions. Users favoring GO interaction specifically highlighted ease of use, with one remarking, “I prefer the GO condition compared to the VG condition,” citing fewer errors. Another shared, “It’s very intuitive to just do the pinching gestures and move your hands.” A user also mentioned, “The task seems to be interesting and refreshing to me,” reinforcing the engaging nature of the system.
Challenges with control and movement accuracy. A prominent theme was the difficulty participants experienced in manipulating the robotic arm accurately. One participant remarked, “It’s not easy to control the robotic arm’s movement accurately,” suggesting improved axis coordination. Another stated, “The task seems to be interesting, but in reality, it was a bit difficult to fully control the robot rotation angles.” Users noted inconsistencies in responsiveness, with one commenting, “Sometimes it was unclear how fast you could move before going into the interruption. The feeling of going into slow movements can make it less efficient.” Issues with sensitivity were also identified: “I think it’s too sensitive and it captures some instability of my hands,” reflecting the need for improved control mapping and clearer visual or tactile cues.
System responsiveness and feedback. Users generally reported that the system was responsive to interaction inputs, with the virtual counterpart reacting immediately upon correct gesture initiation, followed promptly by the physical robot. However, occasional instability – particularly in voice command recognition – was noted, highlighting the need for improved feedback mechanisms and greater system robustness.
V-B2 Voice/Gesture/Multimodal Interaction
@@
Multimodal interaction fluctuated among evaluation. When combining voice and gestures, some articipants experienced increased cognitive load and complexity. Some appreciated the multimodal interaction’s efficiency, especially when both modalities functioned seamlessly. However, participants frequently found switching between voice and gesture inputs unnatural or cumbersome, indicating a need for better integration and smoother transitions. One participant explicitly mentioned, “It was confusing because you have to let it go while you speak and then go back into grabbing cubes.”
@@ V-B3 Suggestions for Improvement
Smoother modality transitions and more fluid robot control. One participant noted challenges in switching between voice and gesture inputs, suggesting that transitions should feel more natural and uninterrupted. Additionally, several users highlighted the need for improved control mapping between the virtual and physical robots. This could be addressed through immediate and clear visual or tactile cues, potentially enhancing overall usability, interaction efficiency, and user awareness.
V-C Analysis Between Roboticist vs. Non-Roboticist
@@
In addition, we explored the differences between roboticists and non-roboticists from our participant pool. Statistical analysis was conducted using mixed methods after the normality tests.
V-C1 Performance
@@
The results indicate that roboticists significantly outperformed non-roboticists in TCT when using merely gesture interaction, completing the task in 194.64 seconds compared to 230.73 seconds for non-roboticists ($p<0.05$) (Fig. 9). However, when voice command was introduced, this difference disappeared, with both groups averaging nearly 300 seconds with no significance. This suggests that GO required a level of expertise that gave roboticists an advantage, while VG effectively equalized performance, making the system more accessible to all users. Despite this, no significant differences were found in other three metrics between the two groups. Similarly, roboticists surpassed non-roboticists obviously in TCT within GO, while both had approximate TCT in VG (Fig. 9), indicating that the task completion was not necessarily dependent on prior robotics experience in multimodal interaction. Regarding unsuccessful attempts, roboticists had an obvious lower mean value in GO, especially in VG, most roboticist had less unsuccessful attempts with the help of voice command interaction (Fig. 9). An interesting trend was observed in the number of interruptions in the VG condition. Roboticists exhibited more interruptions than non-roboticists ($M=5.67\&M=3.94$) (Fig. 9), though this result was not significant. This could suggest that roboticists found the voice commands less efficient or more intrusive, leading to disruptions in their task flow. Non-roboticists, on the other hand, may have benefited from voice assistance, which possibly helped reduce execution errors, leading to fewer system interruptions. This suggests that while roboticists were faster, the overall execution errors and precision challenges were shared across all users, likely influenced more by the interface itself than by individual expertise.
V-C2 UX and Usability
@@
For UEQ-S, both pragmatic and hedonic quality scores showed no significant differences between the two groups (Fig. 10). Roboticists and non-roboticists performed similarly in pragmatic quality under the GO condition, suggesting that the system’s usability was accessible regardless of prior expertise. Notably, non-roboticists reported higher enjoyment and engagement in hedonic quality compared to roboticists. VG consistently yielded slightly higher mean and median values for both quality dimensions than GO, especially for non-roboticists. This indicates that the inclusion of voice interaction may have provided a more intuitive and accessible interface for users unfamiliar with robotic technologies.
@@
NASA TLX revealed more nuanced but functionally meaningful differences. Although no significance emerged, mental demand and effort scores were visibly higher for non-roboticists across both conditions (Fig. 10). This implies that controlling the robot via gesture and voice may have posed greater cognitive challenges for less experienced users. Interestingly, with the addition of voice commands, non-roboticists showed smaller differences from roboticists in temporal demand, performance, and frustration levels, effectively narrowing the workload gap. Notably, frustration levels were marginally higher among roboticists in both conditions, possibly due to their more critical evaluation standards or elevated expectations for system responsiveness and precision based on prior experience.
@@
The most prominent distinctions between the two groups appeared in the Customized Metrics. While statistical significance was not achieved, several sub-metrics revealed clear group differences. Under the GO condition, non-roboticists rated ease of use lower than roboticists, as reflected in both the median values and distribution ranges. However, for all other sub-metrics, non-roboticists gave higher ratings. In the VG condition, non-roboticists outperformed roboticists across all sub-metrics, including a notable improvement in ease of use. This suggests that voice command integration significantly enhanced the subjective experience for novice users, making the system feel more intuitive, controllable, and engaging. On the other hand, the relatively lower ratings from roboticists may reflect their more demanding performance expectations and the importance they place on contextual efficiency when evaluating new systems.
V-C3 Summary of Qualitative Feedback

### v2

@@ 4.4 Measurements
To evaluate the impact of the multimodal interaction, we employed a combination of performance metrics and UX and usability metrics. Performance was measured using four key indicators: number of cubes, unsuccessful attempts, task completion time (TCT, the maximum is 300s), and number of interruptions, providing insights into task efficiency, accuracy, and user control. For UX and usability assessment, we selected UEQ-S (pragmatic quality and hedonic quality, in [-3, +3 Likert scale], with each consisting of four metrics) [schrepp2017design, pick2016design, lin2024usability], a shortened version of UEQ, to balance comprehensive evaluation with experiment duration constraints. Additionally, NASA TLX was utilized to measure the perceived workload in 7-point Likert scale [shayesteh2021investigating, polvi2017handheld]. To further explore usability aspects, we included six customized questions inspired by previous works [zhao2024glanxr, chang2024perceived, lin2024usability], focusing on system usability and user perceptions, which were also measured using 7-point Likert scale for unifying the process as well as better quantifying the evaluation. The complete hierarchy and definitions of self-developed metrics are detailed in Tab. 4.3. We implemented Shapiro–Wilk normality test of each metric and then applied the corresponding statistical analysis.
5 Results
5.1 Quantitative Analysis
5.1.1 Performance Metrics
@@
Number of Cubes. This metric was significantly higher in the GO condition ($M=4.6,SD=2.19$) compared to VG by paired-t test ($M=3.6,SD=1.79,p<0.05$, CohensCohen’s $d=0.60$) with normality ($p>0.05$). The result indicates that incorporating voice commands may have introduced additional cognitive load or inefficiencies, leading to a lower number of successfully moved cubes. The observed variability in performance as shown in Fig. 5(a) suggests differential adaptation to multimodal interaction among participants.
@@
Unsuccessful Attempts. Normality was obtained ($p>0.05$) and paired t-test revels no significant difference was found between the two conditions (CohensCohen’s d=0.12d=-0.12), as shown in Fig. 5(b),. The mean number of failed attempts was slightly higher in VG ($M=7.57,SD=2.41$) than in GO ($M=7.13,SD=3.06$), but this difference was not statistically significant. This implies that error rates were likely influenced by factors such as individual skill levels or task complexity, rather than the interaction modality.
@@
TCT. Wilcoxon signed-rank test was used since normality was not shown in TCT. Given the 300s’ time limit, most participants in VG reached the maximal time, whereas the GO group had a lower TCT ($M=261s$), achieving successful completion more frequently ($p<0.01$, $r=0.19$) (Fig. 5(c)). This finding indicates that gesture-only interaction facilitated more efficient task execution, likely due to fewer errors and quicker responses compared to the multimodal VG condition.
@@
Interruptions. This metric showed normality ($p>0.05$) while the paired t-test displays it happened slightly more frequent in the GO condition ($M=5.95,SD=4.30$) than in VG (Fig. 5(d)). However, this difference was not statistically significant (CohensCohen’s $d=0.22$). The presence of outliers in GO suggests that some users encountered more issues without voice commands, hinting at a possible but inconsistent advantage of multimodal input in reducing interaction errors.
@@
5.1.2 UX and Usability Metrics
UEQ-S. An item-wise analysis of UEQ-S revealed that GO tended to exceed the values across most variables compared to VG (Fig. 6(a)). Despite considerable overlap in distributions, certain items – such as Boring-Exciting – showed clearer separation. Notably, the Inefficient–Efficient scale favored GO significantly (run by paired-t test; $p<0.01$, CohensCohen’s $d=0.23$).
@@ 5.1.2 UX and Usability Metrics
For conventional analysis, pragmatic quality revealed a highly significant difference by Wilcoxon signed-rank test ($p<0.001$, $r=0.35$), with GO ($M=5.19,SD=1.52$) rated higher than VG ($M=4.76,SD=1.64$) (Fig. 6(d)). This suggests that gesture-only interaction was perceived as more efficient, predictable, and user-friendly. Gestures, being direct and immediate, likely contributed to this perception, whereas voice commands may have introduced recognition delays or increased cognitive load. Paired t-test showed hedonic quality differed insignificantly between conditions (CohensCohen’s $d=1.01$), with GO ($M=5.56,SD=1.10$) and VG ($M=5.47,SD=1.14$), which were rated similarly (Fig. 6(e)). This suggests both modalities were comparably engaging and enjoyable. The multimodal novelty of VG did not significantly enhanced user enjoyment to some extent, implying that gesture-based input alone can provide intuitiveness for such experiences. Paired t-test also showed the overall quality significantly favored GO ($M=5.38,SD=1.34$) over VG ($M=5.12,SD=1.46$) ($p<0.001$, CohensCohen’s $d=0.71$) (Fig. 6(f)). This reflects the strong impact of GO’s pragmatic advantages, as hedonic differences were negligible. The results suggest that while voice input may offer contextual benefits, it does not inherently improve UX when combined with gestures. On the contrary, it may introduce complexity and inefficiencies that detract from perceived usability.
@@
NASA TLX. All scales underwent Wilcoxon signed-rank tests since normality was not obtained in any scale. As Fig. 6(b) shows, the VG condition recorded higher scores across most workload categories, contrary to expectations that multimodality would trigger lower scores indicating reduced workload. Notably, VG exhibited more compressed distributions in certain dimensions, pointing to more consistent user perceptions. Significant differences were found in physical demand by paired t-test ($p<0.005$, $r=0.51$), with GO perceived as more physically demanding. This suggests that integrating voice commands alleviates some physical effort. For Perceived Performance, VG was rated significantly higher than GO ($p<0.005$, $r=0.42$), implying that participants felt more successful in task execution when only voice interaction was available. Although mental demand showed no significant difference ($r=0.26$) indicates a slight increase in mental demand for GO. No significances were observed for Temporal Demand, Effort, and Frustration, suggesting similar perceptions of time pressure, exertion, and frustration across both conditions.
Customized Metrics. Compared to standard instruments, customized metrics displayed greater variability (Fig. 6(c)). Wilcoxon signed-rank tests were conducted due to lack of normality in each metric. Some GO metrics had higher medians but wider spread, while VG values tended to be lower and more consistent, indicating possible usability concerns. Significant differences were observed in perceived effectiveness ($p<0.005$, $r=0.33$), perceived intent to use ($p<0.005$, $r=0.42$), and satisfaction ($p<0.005$, $r=0.51$), with GO outperforming VG in all three. This suggests that participants found gesture-only interaction more effective, easier to use, and more satisfying overall. However, the broad variance indicates individual user preferences played a substantial role, especially in other customized metrics not reaching significant differences.
5.2 Qualitative Analysis
We now present the qualitative findings derived from the post-study interviews to gain deeper insight into participants’ experiences and perceptions, obtained from a thematic analysis. One author led the coding process and theme development using an affinity diagram approach, facilitated through a shared Miro board for collaborative access. Two additional authors refined and revised the emerging themes with the assistance of ChatGPT 4o. Finally, the themes were finalized through a participatory discussion among all authors. The main themes with hierarchical details are distilled in Fig. 8.
@@ 5.2 Qualitative Analysis
5.2.1 Overall Experience
Positive impression on intuitiveness & novelty. Participants found the system novel and engaging, appreciating its innovative interaction methods. One participant stated, “I feel like it’s a very novel experience. You can interact with virtual stuff, like the robot model, as well as controlling the real robot. It’s very interesting.” Another noted, “the whole experience is quite nice, everything is very smooth,” underlining the potential for intuitive interactions. Users favoring GO interaction specifically highlighted ease of use, with one remarking, “I prefer the GO condition compared to the VG condition,” citing fewer errors. Another shared, “It’s very intuitive to just do the pinching gestures and move your hands.” A user also mentioned, “The task seems to be interesting and refreshing to me,” reinforcing the engaging nature of the system.
Challenges with control and movement accuracy. While participants appreciated the system’s capabilities, a common theme was that achieving precise manipulation required effort. One participant remarked, “It’s not easy to control the robotic arm’s movement accurately,” suggesting that clearer axis coordination could further assist users. Another stated, “The task seems to be interesting, but in reality, it was a bit difficult to fully control the robot rotation angles.” Users occasionally noted inconsistencies in responsiveness, with one commenting, “Sometimes it was unclear how fast you could move before going into the interruption. The feeling of going into slow movements can make it less efficient.” Sensitivity was also flagged: “I think it’s too sensitive and it captures some instability of my hands,” indicating an opportunity to refine control mapping and provide clearer visual or tactile cues.
System responsiveness and feedback. Users generally reported that the system was responsive to interaction inputs, with the virtual counterpart reacting immediately upon correct gesture initiation, followed promptly by the physical robot, which was positively noted. At the same time, occasional instability – particularly in voice command recognition – was observed, highlighting opportunities to strengthen feedback mechanisms and overall system robustness.
5.2.2 Interaction Modalities
@@
Multimodal interaction fluctuated among evaluation. When combining voice and gestures, some articipants experienced increased cognitive load and complexity. Some appreciated the multimodal interaction’s efficiency, especially when both modalities functioned seamlessly. Some participants found switching between voice and gesture inputs was sometimes unnatural, indicating a need for better integration and smoother transitions. One participant explicitly mentioned, “It was a bit confusing because you have to let it go while you speak and then go back into grabbing cubes.”
@@ 5.2.3 Suggestions for Improvement
Smoother modality transition and more fluid robotic control. One participant noted challenges in switching between voice and gesture inputs, suggesting that transitions should feel more natural and uninterrupted. Additionally, several users highlighted the need for improved control mapping between the virtual and physical robots. This could be addressed through immediate and clear visual or tactile cues, potentially enhancing overall usability, interaction efficiency, and user awareness.
5.3 Analysis Between Roboticist vs. Non-Roboticist
@@
In addition, we explored the differences between roboticists and non-roboticists from our participant pool. Statistical analysis was conducted after the normality tests.
5.3.1 Performance
@@
The results indicate that roboticists significantly outperformed non-roboticists in TCT when using merely gesture interaction, completing the task in 194.64 seconds compared to 230.73 seconds for non-roboticists (independent t-test; $p<0.05$ with a large effect size $Cohen^{\prime}s$ $d=0.926$) (Fig. 8(a)). However, when voice command was introduced, this difference disappeared, with both groups averaging nearly 300 seconds with no significance. This suggests that GO required a level of expertise that gave roboticists an advantage, while VG effectively equalized performance, making the system more accessible to all users. Despite this, no significant differences were found in other three metrics between the two groups. Similarly, roboticists surpassed non-roboticists obviously in TCT within GO, while both had approximate TCT in VG (Fig. 8(c)), indicating that the task completion was not necessarily dependent on prior robotics experience in multimodal interaction. Regarding unsuccessful attempts, roboticists had an obvious lower mean value in GO, especially in VG, most roboticist had less unsuccessful attempts with the help of voice command interaction (Fig. 8(b)). An interesting trend was observed in the number of interruptions in the VG condition. Roboticists exhibited more interruptions than non-roboticists (M=5.67&M=3.94M=5.67\&M=3.94) (Fig. 8(d)), though this result was not significant via Mann–Whitney U test ($r=0.24$). This could suggest that roboticists found the voice commands less efficient or more intrusive, leading to disruptions in their task flow. Non-roboticists, on the other hand, may have benefited from voice assistance, which possibly helped reduce execution errors, leading to fewer system interruptions. This suggests that while roboticists were faster, the overall execution errors and precision challenges were shared across all users, likely influenced more by the interface itself than by individual expertise.
5.3.2 UX and Usability
@@
For UEQ-S, both pragmatic and hedonic quality scores showed no significance between the two groups across the two conditions after performing independent t-tests (Fig. 9(a)). Roboticists and non-roboticists performed similarly in pragmatic quality under the GO condition, suggesting that the system’s usability was accessible regardless of prior expertise. Notably, non-roboticists reported higher enjoyment and engagement in hedonic quality compared to roboticists. VG consistently yielded slightly higher mean and median values for both quality dimensions than GO, especially for non-roboticists. This indicates that the inclusion of voice interaction may have provided a more intuitive and accessible interface for users unfamiliar with robotic technologies.
@@
NASA TLX revealed more nuanced but functionally meaningful differences. Although no significance emerged (normality gained, independent t-tests executed), mental demand and effort scores were visibly higher for non-roboticists across both conditions (Fig. 9(b)). This implies that controlling the robot via gesture and voice may have posed greater cognitive challenges for less experienced users. Interestingly, with the addition of voice commands, non-roboticists showed smaller differences from roboticists in temporal demand, performance, and frustration levels, effectively narrowing the workload gap. Notably, frustration levels were marginally higher among roboticists in both conditions, possibly due to their more critical evaluation standards or elevated expectations for system responsiveness and precision based on prior experience.
@@
The most prominent distinctions between the two groups appeared in the Customized Metrics. While statistical significance was not achieved by independent t-tests, several sub-metrics revealed clear group differences. Under the GO condition, non-roboticists rated ease of use lower than roboticists, as reflected in both the median values and distribution ranges. However, for all other sub-metrics, non-roboticists gave higher ratings. In the VG condition, non-roboticists outperformed roboticists across all sub-metrics, including a notable improvement in ease of use. This suggests that voice command integration significantly enhanced the subjective experience for novice users, making the system feel more intuitive, controllable, and engaging. On the other hand, the relatively lower ratings from roboticists may reflect their more demanding performance expectations and the importance they place on contextual efficiency when evaluating new systems.
5.3.3 Summary of Qualitative Feedback